State whether the following are true or false. Justify your answer. 
 
The value of sin θ increases as θ increases.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The answer is True.

Justification:

According to the values obtained as per the unit circle,
the values of sin are:
sin 0° = 0

sin 30° = 1/2 sin 45° = 1/√2

sin 60° =√3/2

sin 90° = 1
Thus the value of sin θ increases as θ increases.
Note: This answer is in accordance with the value of θ belongs to Ist quadrant only.

Hence, the statement is true

Thanks